Determine the root(s), if they exist, of the quadratic functions:
a. ?(?)=?2−6?+15
b. ?(?)=2?2+3?−2
c. ?(?)=4?2+4?+1
Answer #1

a) کا 23 - 22 - : ( f - To find roots :: رf - گرد ۔ 2 ہلا -b i 12-4ac 2 د را و 2 -ة - : b c=15 رور درج ہے۔ درہ ) + (و۔ ۔ ,ه 2- Зt Jas -3-25 , ас, (Т ч 1 а g, น 2 с) fcn) = 402 tunti +12 +97t, - о агч, b=4, с = 1 -b I √b2-4ac Э За -, t 1 -4 Сус) 2C4)